Minh is 16. His parents are both the same age. The three of them have lived a total of 100. How old are his parents?
Found 2 solutions by robertb, mananth:
Answer by robertb(5830)

16 + x + x = 100 ==> 16 + 2x = 100 ==> 2x = 84 ==> x = 42, the age of each of his parents.

Answer by mananth(16946)

Minh is 16. His parents are both the same age. The three of them have lived a total of 100. How old are his parents?
...
minh 16
father x
mother x
x+x+16 = 100
2x+16=100
add -16 to both sides
2x=100-16
2x= 84
/2
x= 42 parent's age


m.ananth@hotmail.ca